Solve the following recurrences using substitution. (n)T(n 2)3n + 4,for all n 2 3. G iven T(1) = 1, and T(2) 6
Answer #1


(T/M-2) + 3n44 : m=2 n33 T(n) = T (m2) + on the T(n-2) = T (m-4) + 3(n-1) + 4 T(n-4) = T (16) + 3 (-4) +4 substituting T(m) =I Given So, if T(m) T (2) = 6 nok = 2 → k = n-2 from im a T (2) + 3nx T (2) + 3nx n-2 - 2 x 1-2 & 3 + 4x n. 2 + 0/n) - 0(m)